# Iceberg Sensemaking: A Process Model for Critical Data Analysis

### v3

@@ 5.3 Data as a Schematic Artifact
Data is a designed artifact that arises from choices and acts of interpretation made by those who gather it. We assert that you cannot simply lay schemas on top of data; rather, you need schemas to create data in the first place. Every dataset presupposes a schema that gives it structure and meaning. This principle guides our depiction of data as the tip of the schematic iceberg: data is the visible outgrowth of a schematic mass that is largely concealed below the waterline. To overlook the primacy of schemas leads to the illusion that data lives a life apart from the human beings that created it.
None of the three major sensemaking models we analyze adhere to this principle. Every previous sensemaking model takes as a starting point the idea that data is a foundational input, fundamentally unschematized in its raw form. They treat schemas as the ordering frame that we superimpose upon neutral facts in order to bring sense to them. The mindset of treating data as property, where the sensemaking process is intended to make something useful from it, is particularly entrenched in the logic of governments and corporations producing and using datasets. We note that the first sensemaking models emerged from the corporate context of a research lab [21], and were later articulated in the governmental context of intelligence analysis [29].
The mindset of treating data as truth—as raw measurement of objective reality—is very common in the sciences, so a model that sensemaking ”uses data to draw conclusions about the world” [23] feels very natural. The decidedly cognitivist approach of Grolemund & Wickham [22] treats objectivity as the goal, and frames the subjectivity of sensemaking as a flaw to be ameliorated through the data analysis process. In contrast, treating data as a schematic artifact frames data gathering as an intrinsically subjective process. Our model guides analysts on how to confront the entanglement of data with their own contextual subjectivity, and that of data gatherers, rather than attempting to ignore or eliminate this subjectivity.
Nevertheless, we do recognize the utility of the positivist mindset in specific contexts, and do not seek to abandon it entirely: the tip of the iceberg is the place of congruence between our model and previous ones. The difference with the Iceberg model is an expansion of scope that also attends to what happens during the process of gathering the data, before moving on to draw conclusions. Our model insists on the articulation of data's tacit schematization, even if an otherwise positivist stance is responsibly adopted in the sensemaking process.
@@ 5.4 Schematic Multiplicity
While Grolemund & Wickham's [22] approach to data analysis is neatly functional and intuitively powerful, they specifically acknowledge that it suffers from a problem: the potential for an analyst to retain a single false schema throughout the sensemaking process. Our model combats this tendency with the prescriptive guidance to explicitly consider more than one schema.
A single-schema approach to sensemaking lies at the root of many problems in conventional models of data analysis. We argue that these problems can be ameliorated by attending to one's tacitly held schemas up front, explicitly pursuing schematic multiplicity by considering alternative schemas, and interrogating these schemas throughout the sensemaking process.
@@ 6.1 Noticing Uncollected Data
Consider the case of data-driven policing, in which the messy and biased data produced by law enforcement and the courts often reinforce the policing of race and poverty (O’Neil 2016). A skeptic such as the legal scholar Alec Karakatsanis [79], who wants to change the public discourse about how we define criminality, makes data-driven arguments to shine light on contradictions and omissions in how the people involved in law enforcement label and prosecute crimes at different levels of society. One of his core arguments is that only some laws, for some people, are enforced.
6.1.1 Emphasizing Lower-Class Crime
Karakatsanis has illustrated fundamental bias in the US justice system’s pursuit, enforcement, and sentencing of crimes typically committed by different socioeconomic classes [79]. Law enforcement data systematically inflates overpoliced, lower-class crimes such as shoplifting, while also deflating numbers for upper-class crimes such as wage theft that the police are less likely to pursue in their daily work. Using such flawed datasets without attending to this political dimension amounts to a failure of the data sensemaking process.
Consider a scenario in which a civic official uses raw police data to make a straightforward assessment of crime trends. In the Add phase, they acquire a dataset directly from the local police department showing crime trends over the last six months. The official incorporates the explicit schema of this dataset while also absorbing tacit schemas from the dataset itself as well as from their own experience.
@@ 6.1.1 Emphasizing Lower-Class Crime
In the Refine phase, this official does not follow the guidance of the Iceberg Sensemaking model to consider the role of power in the construction of their dataset. As such, they pass over the articulate phase without surfacing elements of the tacit schema that might otherwise refine the naive explicit schema they are actively working with. The update phase may include some simple data wrangling, but otherwise leaves the original dataset mostly untouched. Finally, the official formulates findings based on a dataset, and these findings continue to reflect the general bias of criminal justice data rooted in the over-policing of racialized and lower-class crimes.
@@ 6.2 Learning to Wrangle Data
In addition to the broad implications of the Iceberg Model for how we educate data scientists, it would be immediately effective for recasting lessons on data wrangling [85, 86] as exercises in scrutinizing the construction and limitations of datasets. This scenario follows an illustrative anecdote recounted by boyd [87] in which she encourages data science students to critically explore a dataset before pulling answers from it. Here, the teacher distributes a dataset of police encounters in New York, asking students to find the average age of people apprehended under the city’s controversial and highly racialized Stop and Frisk policy. In the course of the lesson, the students first treat the data as ground truth and report a figure skewed by the messiness of the dataset, then update the data after realizing its messy nature. They finally consider aspects of the data’s construction that might lead to recognizing deeper limitations.
@@ 6.2.1 Getting a Quick Answer
In the Add phase, students first acquire the dataset from their teacher and import it into their analysis software. Because the dataset was gathered and published by the New York Police Department (NYPD), the students have incorporated the explicit schema of concepts, categories, and demographics used by the NYPD while building this dataset. In addition, they have incorporated a range of tacit schemas, including some of the omissions and biases of official data discussed in earlier scenarios. Each student also has absorbed tacit schemas based on their implicit assumptions and past experiences.
In the Check phase, the students begin by interpreting the dataset using only its explicit schema, treating it as an objective basis to calculate the average age of people stopped and frisked by the NYPD. The answer yielded by the dataset is 37. The students quickly evaluate this figure against their tacit schemas and find the number reasonable.
In the Refine phase, the students initially skip past consideration of power and do not articulate tacit schemas that may call it into question. They do not update the dataset before formulating findings based on the raw data, so they reach a straightford answer to the teacher’s question: the students say the average age for offenders recorded in this dataset is 37. The teacher acknowledges that this answer follows directly from the data, but cautions against stopping analysis too soon and encourages the class to loop back and take a closer look.
@@ 6.2.2 Looking Closer at the Data
Returning to the Add phase to start a second loop of analysis, the teacher presents a new tacit schema for students to absorb, pointing out that extensive media coverage of the stop-and-frisk policy found that teenagers were usually targeted, making 37 a rather high figure for an average. She also presents an explicit schema to incorporate into their analysis plan, suggesting that the students should look at the distribution of ages instead of just the mean. The students then jump directly to the Refine phase to articulate this new idea in terms of an explicit schema about the expected age ranges for the targets of this police action.
@@
Although a basic data wrangling lesson might stop there, the instructor also encourages the students to undertake a fourth analysis cycle, with a self-loop back to Refine, to consider power from a critical perspective in the form of incentives for the people gathering the data. Why were so many ages recorded with the placeholder value? The underlying motivation of police in conducting stop and frisk is to exercise disciplinary power in this domain of surveillance and punishment, so the accuracy of the demographic data may be a lesser concern for them. The police may not choose to devote resources to providing clean data if there is no accountability for ensuring its accuracy, such as a structural oversight process that ties senior management bonuses to dataset quality metrics. They may even benefit from obfuscating their activities to reduce public scrutiny. The teacher then invites students to investigate the plausibility of other distributions in the dataset, to potentially surface more limitations before formulating their ultimate findings.
@@ 6.3 Downplaying Inconvenient Data
The Iceberg Sensemaking model is also useful in focusing attention toward the ways that authority figures have interests that may direct the sensemaking process toward their own preferred conclusions. In some data analysis scenarios, the chain of command may introduce pressure to produce certain data analysis results. Consider the case of an analyst at an investment bank who is asked to evaluate a risk model for the bank's investment strategy. An aggressive strategy that minimizes cash reserves would result in very high yields in the short run for the bank, which would result in a multi-million dollar bonus for their manager. However, it presents substantial risks for the bank's liquidity and stability in the long run. The analyst in the following scenario is under considerable pressure from their manager to conclude that the high-yield path is the best approach.
@@ 6.3.1 Acquiescing to Pressure
In the Add phase, the analyst acquires datasets based on the bank's standard risk model. They incorporate both the explicit schemas for these datasets and the tacit schemas that give these datasets context. They also absorb the tacit schema of their manager, who is highly motivated to approve of the strategy that leads to personal enrichment.
@@
In the Refine phase, they choose to change a key assumption in their model, and re-generate a new dataset. In their second analysis loop, returning to the Check phase, the warning disappears. They proceed to the Refine phase and formulate findings that the high-risk strategy is sound. The bank profits immensely, and the manager reaps their millions. However, one year later, the bank fails, because the initial warning was in fact correct. This story of motivated reasoning essentially describes what happened with the Silicon Valley bank failure of 2023 (Gilbert et al. 2023).
@@ 6.3.2 Questioning Authority
We now consider how adhering to the Iceberg model could have changed the sensemaking outcome in this scenario. The Add phase proceeds as above. However, in the Check phase, the analyst takes the warning arising from the dataset seriously and evaluates whether their schemas harbor faulty assumptions. In the Refine phase, they consider the role of interpersonal power in their assumptions, and realize that the pressure from their manager to green-light the pursuit of short-term profits should not lead them to dismiss the warning too hastily. Faced with a mismatch between dataset and explicit schema, they could either interpret the dataset as incorrect, or evaluate whether the explicit schema may be incorrect. Informed by their consideration of power, they articulate an explicit schema recognizing that very real risks may threaten the long-term viability of the bank itself.
@@
A skeptic might say “this is just the result of bad analysis.” In fact, that’s the point: other sensemaking models would fail to account for the role of power in the workplace and its effect on sensemaking. The absence of political considerations in other sensemaking models means these models are unable to explain cases where power knocks data analysis off the rails.
@@ 6.4 Measuring With Sensors
We finally consider the question of sensemaking with datasets that arise from measurements taken with sensors. At first glance, it may be tempting to assume that all such contexts are prime territory for a positivist framing, where interpretivism is an unnecessary diversion of attention. However, the Iceberg Model has explanatory value in showing how the sensemaking process can fail when bad data is certified as official and reliable while alternative data sources are ignored. Our model also has prescriptive value in showing how to avoid these failures.
@@ 6.4.2 Recognizing a Water Crisis
Imagine a new analyst who takes over water quality measurement in a city like Flint, shortly after governmental officials admit to problems in the water testing regime. In the Add phase, the analyst acquires data by taking samples of the city’s drinking water using the procedure established by their predecessor. They incorporate existing explicit schemas, including directives on where and how often to gather and test the water samples. At the same time, the limitations of this dataset are also incorporated as an implicit schema because this is essential context that has yet to be articulated. From our privileged perch as post-hoc observers of a water crisis, we recognize that the full story behind the data has yet to be revealed because it lies implicit in a tacit schema. The tacit schema also includes the analyst’s personal knowledge, such as news reports of city residents forced to rely entirely on bottled water because their tap water is undrinkable. This knowledge is absorbed into the analysis process as a tacit schema.
In the Check phase, the analyst interprets the given data and sees similar figures to those reported by their predecessor. The fact that the data show perfectly safe toxin levels leads them to evaluate their tacit schema and wonder whether reports of unsafe water are exaggerated, or if the dataset itself is flawed. Here, the value of schematic multiplicity is evident in the serious consideration of different hypotheses that may require new data and alternative forms of analysis.
@@
Beyond this specific scenario, it is worth noting that political complications frequently emerge in the broader sphere of public health data. The Covid-19 pandemic frequently showcased how debates over the limitations of various datasets and the framing of scientific results amounted to sensemaking complications in which multiple schemas were set at odds, often resulting in contentious reexamination of the sensemaking process and how it led to one conclusion or another [89].
@@ 6.4.3 Quantifying Rainfall
Finally, consider a fairly straightforward data analysis scenario: rainfall sensors used by a government scientist to assess drought trends in their state. The sensors are placed at convenient locations to facilitate periodic calibration, and the data is gathered remotely, stored in a spreadsheet, and later analyzed to look for short-term changes and long-term trends.
@@
In this scenario, it is fine to work in the basic mode of positivism, treating measurements as a straightforward proxy for the phenomenon under study, while focusing attention and scrutiny mainly on sensor calibration, the soundness of your model, and accurate presentation of statistical findings. In cases like this, positivism is expedient and uncomplicated, granting speed and simplicity to the sensemaking process. The same is true in a case like air-traffic control, where a positivistic sense of certainty is generally preferable to critical doubting of the data. Our basic position is that data analysts must perform some due diligence in the examination of power before embracing a positivist stance, rather than simply assuming that a straightforward positivist approach to their subject is warranted. The Iceberg Sensemaking model thus accommodates positivism as a matter of epistemic pluralism, recognizing it as productively compatible within a broader interpretivist framework designed to check for socio-political complications in the sensemaking process.
For this reason, previous sensemaking models could produce adequate descriptive and explanatory accounts of the rainfall measurement scenario because the role of tacit schemas and power is largely uncomplicated. Yet such uncomplicated circumstances are not ideal for testing the value of a sensemaking model, which should also explain how false and misleading interpretations emerge in problematic cases. While a positivist approach to sensemaking may have descriptive power in well understood cases, it has poor explanatory value in situations where sensemaking requires context, skepticism, and critique—qualities that characterize the best scientific work, even if these hunches and other insights happen to remain tacit and unarticulated.

### v4

@@ 5.3 Data as a Schematic Artifact
Data is a designed artifact that arises from choices and acts of interpretation made by those who gather it. We assert that you cannot simply lay schemas on top of data; rather, you need schemas in order to create data in the first place. Every dataset presupposes a schema that gives it structure and meaning. This principle guides our depiction of data and explicit schemas as the tip of the tacit schematic iceberg: data is the visible outgrowth of a schematic mass that is largely concealed below the waterline. To overlook the primacy of schemas leads to the illusion that data lives a life apart from the human beings that created it.
None of the three major sensemaking models we analyze adhere to this principle. Every previous sensemaking model takes as a starting point the idea that data is a raw input, fundamentally unschematized in its natural form. This approach leads schemas to be treated as the ordering frame that we impose upon neutral facts to make them useful. The mindset of treating data as a form of material property, in which the sensemaking process is intended to make something useful from raw input, is particularly entrenched in the logic of governments and corporations producing and using datasets. We note that the first sensemaking models emerged from the corporate context of a research lab [21], and were later articulated in the governmental context of intelligence analysis [29].
The mindset of treating data as ground truth—as raw measurement of objective reality—is very common in the sciences, so a model that sensemaking ”uses data to draw conclusions about the world” [23] feels very natural. The decidedly cognitivist approach of Grolemund & Wickham [22] treats objectivity as the goal, and frames the subjectivity of sensemaking as a flaw to be ameliorated through the data analysis process. In contrast, treating data as a schematic artifact frames data gathering as an intrinsically subjective process, even though it may be useful to treat data as objective in certain methodological contexts. Our model guides analysts on how to confront the entanglement of data with their own contextual subjectivity, and that of data gatherers, rather than ignore or attempt to eliminate subjectivity.
Nevertheless, our model does incorporate the utility of the positivist mindset in specific contexts, so we do not wish to abandon it entirely: the sections of the iceberg standing above the water are largely congruent between our model and previous ones. The difference with the Iceberg model is an expansion of scope that also attends to what happens during the process of gathering the data, before moving on to draw conclusions. Our model insists on the articulation of data’s tacit schematization, even if an otherwise positivist stance is responsibly adopted in the sensemaking process. The Iceberg model is a pragmatic effort to synthesize interpretivist approaches within the full scope of existing methodologies, styles of work, and values of visualization as a field.
@@ 5.4 Schematic Multiplicity
While Grolemund & Wickham’s [22] approach to data analysis is neatly functional and intuitively powerful, they specifically acknowledge that it suffers from a problem: the potential for an analyst to retain a single false schema throughout the sensemaking process. Our model targets this tendency with prescriptive guidance to explicitly consider more than one schema.
A single-schema approach to sensemaking lies at the root of many problems in conventional models of data analysis. We argue that these problems can be ameliorated by attending to one’s tacitly held schemas up front, explicitly pursuing schematic multiplicity through considering alternative schemas, and interrogating these schemas throughout the sensemaking process.
@@ 6.1 Noticing Uncollected Data
Consider the case of data-driven policing, in which the messy and biased data produced by law enforcement and the courts often reinforce the policing of race and poverty [6]. A skeptic such as the legal scholar Karakatsanis [80], who wants to change the public discourse about how we define criminality, makes data-driven arguments to shine light on contradictions and omissions in how the people involved in law enforcement label and prosecute crimes at different levels of society. One of his core arguments is that only some laws, for some people, are enforced.
6.1.1 Reexamining Lower-Class Crime
Karakatsanis has illustrated fundamental bias in the US justice system’s pursuit, enforcement, and sentencing of crimes typically committed by different socioeconomic classes [80]. Law enforcement data systematically inflates overpoliced, lower-class crimes such as shoplifting, while also deflating numbers for upper-class crimes such as embezzlement and wage theft that the police are less likely to pursue in their daily work. Using such flawed datasets without attending to their political dimensions amounts to a failure of the data sensemaking process.
Consider a scenario in which a civic official uses police data to make a straightforward assessment of crime trends. In the Add phase, they acquire a dataset directly from the local police department showing crime trends over the last six months. The official incorporates the explicit schema of this dataset while also absorbing tacit schemas from the dataset itself as well as from their own experience.
@@ 6.1.1 Reexamining Lower-Class Crime
In the Refine phase, this official does not follow the guidance of the Iceberg Sensemaking model to consider the role of power in the construction of their dataset. As such, they pass over the articulate phase without surfacing elements of the tacit schema that might otherwise refine the naive explicit schema they are actively working with. The update phase may include some simple data wrangling, but otherwise leaves the original dataset mostly untouched. Finally, the official formulates findings based on their dataset, and these findings continue to reflect the general bias of criminal justice data rooted in the over-policing of racialized and lower-class crimes.
@@ 6.2 Learning to Wrangle Data
In addition to the broad implications of the Iceberg Model for how we educate data scientists, this model would be immediately effective for recasting lessons on data wrangling [89, 90] as exercises in scrutinizing the construction and limitations of datasets. The following scenario uses an illustrative classroom anecdote recounted by boyd [91] in which she encourages data science students to critically explore a dataset before pulling answers from it. Here, the teacher distributes a dataset of police encounters in New York, asking students to find the average age of people apprehended under the city’s controversial and highly racialized Stop and Frisk policy. In the course of the lesson, the students first treat the data as ground truth and report a figure skewed by the messiness of the dataset, then update the data after realizing its messy nature. They finally consider aspects of the dataset’s construction that might lead to recognizing deeper limitations.
@@ 6.2.1 Getting a Quick Answer
In the Add phase, students first acquire the dataset from their teacher and import it into their analysis software. Because the dataset was gathered and published by the New York Police Department (NYPD), the students have incorporated the explicit schema of concepts, categories, and demographics used by the NYPD while building this dataset. In addition, they have incorporated a range of tacit schemas, including some of the omissions and biases of official data discussed in earlier scenarios. Each student also has absorbed a set of tacit schemas based on their implicit assumptions and past experience.
In the Check phase, the students begin by interpreting the dataset using only its explicit schema, treating it as an objective basis to calculate the average age of people stopped and frisked by the NYPD. The answer yielded by the dataset is 37. The students quickly evaluate this figure against their unarticulated tacit schemas and find the number reasonable.
In the Refine phase, the students initially skip past consideration of power and do not articulate tacit schemas that may call it into question. They do not update the dataset before formulating findings based on the raw data, so they reach a straightforward answer to the teacher’s question: the students say the average age for offenders recorded in this dataset is 37. The teacher acknowledges that this answer follows directly from the data, but cautions against stopping analysis too soon and encourages the class to loop back and take a closer look.
@@ 6.2.2 Looking Closer at the Data
Returning to the Add phase to start a second loop of analysis, the teacher presents a new tacit schema for students to absorb, pointing out that extensive media coverage of the stop-and-frisk policy found that teenagers were usually targeted, making 37 a rather high figure for an average. She also presents an explicit schema to incorporate into their analysis plan, suggesting that the students should look at the distribution of ages instead of just the mean. The students then jump to the Refine phase, consider power, and articulate this new idea in terms of an explicit schema about the expected age ranges for the targets of this police action.
@@
Although a basic data wrangling lesson might stop there, the instructor also encourages the students to undertake a fourth analysis cycle, with a self-loop back to Refine, to again consider power from a critical perspective in the form of incentives for the people gathering the data. Why were so many ages recorded with the placeholder value? The underlying motivation of police in conducting stop and frisk is to exercise disciplinary power in this domain of surveillance and punishment, so the accuracy of the demographic data may be a lesser concern for them. The police may not choose to devote resources to providing clean data if there is no accountability for ensuring its accuracy, such as an oversight process that ties senior management bonuses to dataset quality metrics. They may even benefit from obfuscating their activities to reduce public scrutiny. The teacher then invites students to investigate the plausibility of other distributions in the dataset, to potentially surface more limitations before formulating their ultimate findings.
@@ 6.3 Downplaying Inconvenient Data
The Iceberg Model is also useful in focusing attention toward the ways that authority figures have interests that may direct the sensemaking process toward their own preferred conclusions. In some data analysis scenarios, the chain of command may introduce pressure to produce certain data analysis results. Consider the case of an analyst at an investment bank who is asked to evaluate a risk model for the bank’s investment strategy. An aggressive strategy that minimizes cash reserves would result in very high yields in the short run for the bank, which would result in a multi-million dollar bonus for their manager. However, it presents substantial risks for the bank’s liquidity and stability in the long run. The analyst in the following scenario is under considerable pressure from their manager to conclude that the high-yield path is the best approach.
@@ 6.3.1 Acquiescing to Pressure
In the Add phase, the analyst acquires datasets based on the bank’s standard risk model. They incorporate both the explicit schemas for these datasets and the tacit schemas that give these datasets context. They also absorb the tacit schema of their manager, who is highly motivated to approve of the strategy that leads to personal enrichment.
@@
In the Refine phase, the analyst chooses to change a key assumption in their model, and re-generate a new dataset. In their second analysis loop, returning to the Check phase, the warning disappears. They proceed to the Refine phase and formulate findings that the high-risk strategy is sound. The bank profits immensely, and the manager reaps their millions. However, one year later, the bank fails, because the initial warning was in fact correct. This story of motivated reasoning essentially describes what happened with the Silicon Valley bank failure of 2023 [92].
@@ 6.3.2 Questioning Authority
We now consider how adhering to the Iceberg model could have changed the sensemaking outcome in this scenario. The Add phase proceeds as above. However, in the Check phase, the analyst takes the warning arising from the dataset seriously and evaluates whether their schemas harbor faulty assumptions. In the Refine phase, they consider the role of interpersonal power in their assumptions, and realize that the pressure from their manager to green-light the pursuit of short-term profits should not lead them to dismiss the warning too hastily. Faced with a mismatch between dataset and explicit schema, they could either interpret the dataset as incorrect or evaluate whether the explicit schema may be incorrect. Informed by their consideration of power, they articulate an explicit schema recognizing that very real risks may threaten the long-term viability of the bank itself.
@@
A skeptic might say “this is just the result of bad analysis”. In fact, that’s the point: other sensemaking models would fail to account for the role of power in the workplace and its effect on sensemaking. Similar intrusions of power on the sensemaking process even occur in the realm of scientific research, where there is considerable pressure to produce positive results. The absence of political considerations in other sensemaking models means these models are unable to explain cases where power knocks data analysis off the rails.
@@ 6.4 Measuring With Sensors
We now consider the question of sensemaking with datasets that arise from measurements taken with sensors. At first glance, it may be tempting to assume that mechanical data gathering is prime territory for a positivist framing, where interpretivism is an unnecessary diversion of attention. However, the Iceberg Model has explanatory value in showing how the sensemaking process can fail when bad data is certified as official and reliable while alternative data sources are ignored. Our model also has prescriptive value in showing how to avoid these failures.
@@ 6.4.2 Recognizing a Water Crisis
Imagine a new analyst who takes over water quality measurement in a city like Flint, shortly after governmental officials admit to problems in the water testing regime. In the Add phase, the analyst acquires data by taking samples of the city’s drinking water using the procedure established by their predecessor. They incorporate existing explicit schemas, including directives on where and how often to gather and test the water samples. At the same time, the limitations of this dataset are also incorporated as an implicit schema because this is essential context that has yet to be articulated. From our privileged perch as post-hoc observers of this water crisis, we recognize that the full story behind the data has yet to be revealed because it lies implicit in a tacit schema. The tacit schema also includes the analyst’s personal knowledge, such as news reports of city residents forced to rely entirely on bottled water because their tap water is undrinkable. This knowledge is absorbed into the analysis process as a tacit schema.
In the Check phase, the analyst interprets the given data and sees similar figures to those reported by their predecessor. The fact that the dataset shows perfectly safe toxin levels leads them to evaluate their tacit schemas and wonder whether reports of unsafe water are exaggerated, or if the dataset itself is flawed. Here, the value of schematic multiplicity is evident in the serious consideration of different hypotheses that may require new data and alternative forms of analysis.
@@
Beyond this specific scenario, it is worth noting that political complications frequently emerge in the broader sphere of public health data. The Covid-19 pandemic frequently showcased how debates over the limitations of various datasets and the framing of scientific results amounted to sensemaking complications in which multiple schemas were set at odds, often resulting in contentious reexamination of the sensemaking process and how it led to one conclusion versus another [94].
@@ 6.4.3 Quantifying Rainfall
Finally, consider a fairly straightforward data analysis scenario: rainfall sensors used by a scientist to assess drought trends in their state. The sensors are placed at convenient locations to facilitate periodic calibration, and the data is gathered remotely, stored in a spreadsheet, and later analyzed to look for short-term changes and long-term trends.
@@
In this scenario, it is fine to work in the basic mode of positivism, treating measurements as a straightforward proxy for the phenomenon under study, while focusing attention and scrutiny mainly on sensor calibration, the soundness of your model, and accurate presentation of statistical findings. In cases like this one, positivism is expedient and uncomplicated, granting speed and simplicity to the sensemaking process. The same is true in a case like air-traffic control, where a positivistic sense of certainty is generally preferable to critical doubting of the data, even though the controller should also be attuned to potential signs that they may need to doubt their instruments. Our basic position is that data analysts must perform some due diligence in the examination of power and the reliability of data before embracing a positivist stance, rather than simply assuming that a straightforward positivist approach to their subject is warranted. The Iceberg Sensemaking model thus accommodates positivism as a matter of epistemic pluralism, recognizing it as productively compatible within a broader interpretivist framework designed to check for socio-political complications and procedural dysfunction in the sensemaking process.
For this reason, previous sensemaking models could produce adequate descriptive and explanatory accounts of the rainfall measurement scenario because the role of tacit schemas and power is largely uncomplicated. Yet such uncomplicated circumstances are not ideal for testing the value of a sensemaking model, which should also explain how false and misleading interpretations emerge in problematic cases. While a positivist approach to sensemaking may have descriptive power in well understood cases, it has poor explanatory value in situations where sensemaking requires context, skepticism, and critique—qualities that truly characterize the best scientific work, even if these hunches [96] and other insights happen to remain tacit and unarticulated.